The perimeter of a room is 66 feet.the ratio of its length to its width is 6:5. you want to tile the floor with 12 inch square tiles. find the length and width of the room and the area of the floor. how many tiles will you need ?the tile cost $1.98 each. what is the total cost to tile the floor ?
Answer by mananth(16946)

Perimeter = 66 feet
ratio of length to width = 6/5
2(6+5) =22 parts
..
Length = 6/22*66= 18
Width = 5/22*66 = 15
..
Area = 18*15 = 270 sq. feet
Area of tile = 1 sq.ft
...
number of tiles = 270 tiles
..
1 tile costs 1.98
270 tiles cost 1.98 * 270
= $ 534.60